• ERAM-46,47,48 (last 3 ERAM) production is on-going
  • status of ERAM-38-45
    • ERAM-38 : at cabling division (6 dead pads)
    • ERAM-39 : scanned OK
    • ERAM-40 : mesh pulsing OK, Glued. scan is on-going.
    • ERAM-41 : scanned OK
    • ERAM-42 : 2-3 pads connected to each other. 10 rms noise spread over the surface (58 pads). To understand before gluing.
    • ERAM-43 : mesh pulsing done, OK (no dead pad). glued, ready for scan.
    • ERAM-44 : mesh pulsing done
    • ERAM-45 : mesh pulsin done
  • 3 stiffener remaining for ERAM-44 & 45. One to be sandblasted. 
  • Production of 6 stiffener (external company) : paused because the O-ring groove is not in the production drawings !
    • Fabrizio is awaiting feedback from henry (IFJPAN)
    • Charlotte will try to get the groove measured by CERN metrology : given today.